40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 3 21:05:23 2022, yuta, Update, BHD, AS path alignment, removing a lot of green stuff 7x
    Reply  Wed May 4 13:09:42 2022, Anchal, Update, BHD, IMC table balanced again, IMC is locking, YARM is locking 
       Reply  Wed May 4 15:08:59 2022, Tega, Update, BHD, XARM alignment to get flashing IFO_XY_ARMS_Flashing.png
          Reply  Wed May 4 18:52:43 2022, Tega, Update, BHD, POP beam too high at POP SM4 on the ITMX chamber snnp_pr2_pr3_itmy_aligment.png
             Reply  Thu May 5 14:46:22 2022, Anchal, Update, BHD, POP beam height lowered, POP_SM4 raised 
                Reply  Thu May 5 17:05:31 2022, Tega, Update, BHD, IMC & X/Y-arm alignment IFO_aligment_Y_locking_on.png
                Reply  Mon May 9 18:43:03 2022, Anchal, Update, BHD, ITMX table layout corrected 
                   Reply  Mon May 9 18:49:05 2022, Tega, Update, BHD, Relocate green TRX and TRY PDs/cams/optics from PSL table to BS table IMG_20220509_184439943.jpgIMG_20220509_184154722.jpg
                      Reply  Tue May 10 15:46:38 2022, JC, Update, BHD, Relocate green TRX and TRY components from PSL table to BS table IMG_0741.jpegIMG_0753.jpeg
                         Reply  Tue May 10 18:25:43 2022, Paco, Update, BHD, Green Transmission path 
                   Reply  Tue May 10 16:31:14 2022, Tega, Update, BHD, ITMX optlev return beam steered to QPD 
       Reply  Fri May 6 13:48:34 2022, Anchal, Update, BHD, WFS loop instability fixed 
    Reply  Thu May 12 19:55:01 2022, Tega, Update, BHD, AS path alignment AS.JPGOMCchamber.JPG
       Reply  Thu May 12 20:11:18 2022, Anchal, Update, BHD, BHDBS Output beams steered out to ITMY table signal-2022-05-12-201844.jpeg
          Reply  Thu May 12 20:24:29 2022, yuta, Update, BHD, POX and POY investigation POY.JPG
       Reply  Mon May 16 13:22:59 2022, yuta, Update, BHD, REFL and AS paths aligned at AP table REFLPOP.JPGPOPAS.JPG
          Reply  Mon May 16 19:14:17 2022, Anchal, Update, BHD, Camera set on AS path and BHDBS output path 
             Reply  Tue May 17 18:43:38 2022, Anchal, Update, BHD, Placed SRM in ITMY Chamber 
                Reply  Wed May 18 08:30:29 2022, Paco, Update, BHD, SRM OpLev 
                   Reply  Thu May 19 08:51:40 2022, Paco, Update, BHD, SRM OpLev 
                Reply  Wed May 18 09:02:52 2022, Anchal, Update, BHD, WFS1 PD centered 
                Reply  Wed May 18 17:23:15 2022, Anchal, Update, BHD, Placed PRM in BS Chamber 
                   Reply  Thu May 19 18:56:08 2022, yuta, Update, BHD, Current OSEM sensor values with all the suspensions aligned OSEMS.txt
                   Reply  Thu May 19 19:05:59 2022, Anchal, Update, BHD, BS Chamber all work finished, BHD path setup 
                      Reply  Fri May 20 20:03:48 2022, Paco, Update, BHD, ITMY chamber work finished - LO and AS overlapped Screenshot_2022-05-20_20-38-25.pngOSEMs.txt
                         Reply  Mon May 23 13:16:59 2022, Paco, Update, BHD, End of vent - checks OSEMs20220523.txtScreenshot_2022-05-23_13-24-06_IFOaligned_GTRY_still_clipping_but_less.pngPR3position.JPG
                            Reply  Tue May 24 15:21:13 2022, Anchal, Update, BHD, Freeswing tests of new SOS started 
Message ID: 16869     Entry time: Mon May 23 13:16:59 2022     In reply to: 16868     Reply to this: 16872
Author: Paco 
Type: Update 
Category: BHD 
Subject: End of vent - checks 

[Paco, Yuta]

Prep for closing and pump down.

  • Aligned IFO to maximize DC levels.
    • YARM (flashing peak 0.05 with PRM misaligned), XARM (flashing peak 0.06 with PRM misaligned), PRC (PRY flashing -30 @ POPDC, offset -70 and REFL DC 270), SRC (SRY flashing -30 @ POPDC, offset -70), BHD.
  • GTRY clipping
    • We tried moving the alignment of PR3, PR2, ETMY, ITMY to reduce clipping and retain IR flashing. We found it kind of difficult, so we only used the unclipped GTRY temporarily to improve the input YAUX injection after which the YAUX locked. We then restored the clipping in favor of the IR beam alignment.
  • PR3 position
    • PR3 seems to be +1 inch away towards East, nominally placed along North-South, and offset in YAW.
  • Aligned OPLEVs to center at around Mon May 23 13:20:32 2022
  • Snapshot of all cameras in the control room around Mon May 23 13:24:51 2022

[Chub, JC, Jordan, Yuta, Yehonathan, Paco]

Closed in the following order:

  • IMC chamber
  • OMC chamber
  • BS chamber
  • ITMY chamber
  • ITMX chamber

[Yuta, Paco]

After closing the heavy doors, we tried to have GTRY less clipped using PR2, PR3, ITMY and ETMY. During this adventure, we also aligned GRY injection beam by hand. Rotating a waveplate for GRY injection made GRY locking stably at GTRY of ~0.3.

Attachment 1: OSEMs20220523.txt  2 kB  Uploaded Mon May 23 21:11:01 2022  | Hide | Hide all
2022-05-23 13:15:59.531109
 =====BS=====
C1:SUS-BS_ULSEN_OUT16 =   600 
C1:SUS-BS_LLSEN_OUT16 =   575 
C1:SUS-BS_URSEN_OUT16 =   600 
C1:SUS-BS_LRSEN_OUT16 =   635 
C1:SUS-BS_SDSEN_OUT16 =   670 
 =====ITMX=====
C1:SUS-ITMX_ULSEN_OUT16 =   404 
C1:SUS-ITMX_LLSEN_OUT16 =   608 
... 76 more lines ...
Attachment 2: Screenshot_2022-05-23_13-24-06_IFOaligned_GTRY_still_clipping_but_less.png  403 kB  Uploaded Mon May 23 21:11:56 2022  | Hide | Hide all
Screenshot_2022-05-23_13-24-06_IFOaligned_GTRY_still_clipping_but_less.png
Attachment 3: PR3position.JPG  5.195 MB  Uploaded Mon May 23 21:21:43 2022  | Hide | Hide all
PR3position.JPG
ELOG V3.1.3-